Coastal Star Sales Corporation Questions Calculate a mean and a standard deviation for each variable. 2. Set a 95 percent con help with questions 3 and 4
Answer #1

3.

The quantitative variables in the data provided are Age, Years of Experience, Previous Year Sales, Current Year Sales.

So the median, mode, range are being calculated for these four quantitative variables.

At first, the data is to be arranged in ascending order.

Number of observations = 11

So :- Median = [(11+1)/2]th observation = 6th observation

Mode = 11th observation

Range = 11th observation - 1st observation

These are achieved using R-programming as follows :-

age=c(40,60,26,39,64,51,34,62,42,64,27)
years_of_experience=c(7,12,2,1,5,2,1,10,3,5,2)
previous_year_sales=c(412744,1491024,301421,401241,448160,518897,846222,1527124,921174,463399,548011)
current_year_sales=c(411007,1726630,700112,471001,449261,519412,713333,2009041,1030000,422798,422001)
age_median=sort(age)[6]
years_of_experience_median=sort(years_of_experience)[6]
previous_year_sales_median=sort(previous_year_sales)[6]
current_year_sales_median=sort(current_year_sales)[6]
age_mode=sort(age)[11]
years_of_experience_mode=sort(years_of_experience)[11]
previous_year_sales_mode=sort(previous_year_sales)[11]
current_year_sales_mode=sort(current_year_sales)[11]
age_range=sort(age)[11]-sort(age)[1]
years_of_experience_range=sort(years_of_experience)[11]-sort(years_of_experience)[1]
previous_year_sales_range=sort(previous_year_sales)[11]-sort(previous_year_sales)[1]
current_year_sales_range=sort(current_year_sales)[11]-sort(current_year_sales)[1]
age_median;age_mode;age_range
years_of_experience_median;years_of_experience_mode;years_of_experience_range
previous_year_sales_median;previous_year_sales_mode;previous_year_sales_range
current_year_sales_median;current_year_sales_mode;current_year_sales_range

The output :-

age_median = 42 age_mode = 64 age_range = 38
years_of_experience_median = 3 years_of_experience_mode = 12 years_of_experience_range = 11
previous_year_sales_median = 518897 previous_year_sales_mode = 1527124 previous_year_sales_range = 1225703
current_year_sales_median = 519412 current_year_sales_mode = 2009041 current_year_sales_range = 1598034

Median of Age = 42 years Mode of Age = 64 years Range of Age = 38 years

Median of Years of Experience = 3 years   
Mode of Years_of_Experience = 12 years Range of Years of Experience = 11 years

Median of Previous Year Sales = $518897
Mode of Previous Year Sales = $1527124
Range of Previous Year Sales = $1225703

Median of Current Year Sales = $519412
Mode of Current Year Sales = $2009041   
Range of Current Year Sales = $1598034

.

.

\textbf{4.}

Class Tally Marks Frequency
<500000 ||||| 5
500000-1000000 ||| 3
>1000000 ||| 3

Total Frequency = 5+3+3 = 11